OpenStudy (anonymous):

12/1 divided (3b/2) cubed

OpenStudy (shane_b):

Can you use the equation editor or draw it out? Not sure exactly what's being cubed there.

OpenStudy (anonymous):

\[12\div (3b \div2)^{3}\]

OpenStudy (shane_b):

Ok \[\large \frac{12}{(\frac{3b}{2})^3}=>\frac{12}{(\frac{3b^3}{2^3})}=>12*\frac{8}{3b^3}=>\frac{96}{3b^3}=>\frac{32}{b^3}\]

OpenStudy (anonymous):

what about this \[solve for a : 6a+5/3 - 2a+2/2 =8/2\]

OpenStudy (callisto):

Please close the question and start a new post for a new question using the close button.

OpenStudy (anonymous):

different one its solve for a: 6a+5/3 - 2a+2/2 =8/2